RC meeting and Run Plan, Wednesday November 3rd

From SBS wiki
Jump to: navigation, search

Meetings Page

RC Meeting Information

RC Daily Meetings are at 3:15pm, remote connection link: https://bluejeans.com/256205934

Accelerator Status

https://logbooks.jlab.org/entry/3936389

Hall A Status and News

Subsystem Updates

  1. HCAL: N/A
  2. BBCal: still need to order parts for dry air flow to PMT; scatter magnetic field sensors around BBPS to evaluate fringe fields: shall they be removed before taking beam? answer: we don't know about radiation hardness of these, but hey are pretty cheap; they shall be fixed in such a way they won't move around;
  3. Timing Hodoscope: calibrations: time walks, light propagation speeds, etc msotly ready, shall be submitted for pull request tomorrow;
  4. Grinch: need to change FADC mode from 1 to 7; need to take LED data to test;
  5. GEMS: took cosmics last night to check the latency, need to go downstairs to check the LV plug.
  6. Beamline: scalers from SBSVME29 look sensible, unlike LHRS scalers. Most likely it is a software issue, which is investigated by David and Bob.
  7. DAQ: N/A
  8. Replay Analysis: N/A